# VFEFL: Privacy-Preserving Federated Learning against Malicious Clients via Verifiable Functional Encryption

### v2

@@ Abstract
Index Terms:
privacy protection, federated learning, functional enctyption.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Federated learning[1] is an advanced and promising distributed learning paradigm that enables multiple clients to collaboratively train models without sharing local data. It embodies the principles of focused data collection and minimization, thereby mitigating many of the systemic privacy risks and costs inherent in traditional, centralized machine learning approaches. With the enactment of privacy protection and data security regulations in various countries, as well as the increase in public privacy awareness, the application of federated learning has gradually expanded e. g., healthcare, finance, and edge computing.
Although the federated learning framework offers a degree of privacy protection for client data during training, it brings new privacy and security issues [2]. One primary concern is the potential for privacy leakage through the local models uploaded by clients. With the advent and ongoing advancement of attack techniques, such as model inversion attacks[3, 4, 5], submitting local models in plaintext form is no longer regarded as a secure method. Accordingly, obfuscation and encryption mechanisms have been adopted to address this issue. Another threat is attacks from malicious clients. Due to its distributed architecture, federated learning is susceptible to adversarial manipulation by malicious clients. A malicious client may make the accuracy of the global model drop rapidly by uploading malicious models, or destroy the decryption by uploading a malicious decryption key or a wrong ciphertext. The above issues are important for federated learning systems to achieve their core goal (training high quality models while preserving data privacy). Privacy-preserving mechanisms (such as encryption schemes) can potentially lead to difficulties in detection and verification, while detecting and verifying encrypted models of the clients may compromise specific information. Existing solutions [6, 7] cannot be deployed in the basic federated learning architecture (one server and multiple clients), and need help from two non-colluding servers or additional trusted third parties, which introduces additional limitations and consumption.
@@
    We propose a novel DVFE scheme that preserves data privacy and ensures correct encryption and decryption. It supports verification of relations over multidimensional ciphertexts and is suitable to federated learning. Additionally, we formalize its definition, security model, and reduce its security in the proposed security model to well-known complexity assumptions.
@@ II-A Functional Encryption
Functional encryption (FE) [8, 9] is a paradigm supports selective computation on encrypted data, while traditional public key encryption only implements all-or-nothing decryption method. An authority in a FE scheme can generate restricted decryption keys to allow receiver to learn specific functions of the encrypted messages and nothing else. Goldwasser[10] first introduced multi-input functional encryption (MIFE), which is a generalization of functional encryption to the setting of multi-input functions. Michel Abdalla [11] constructed the first MIFE for a non-trivial functionality with polynomial security loss under falsifiable assumptions for a super-constant number of slots.
However, whereas the input can be large, the original definition of FE limits the source of the encrypted data to a single client, which is not compatible with a wide range of real-world scenarios where data needs to be collected from multiple sources. Multi-client functional encryption (MCFE)[12] was proposed to addresses this issue. Further, decentralized multi-client functional encryption (DMCFE)[12] eliminated reliance on a trusted party. In the DMCFE scheme instantiated over an asymmetric pairing group, each client independently samples $s_{i}\leftarrow\mathbb{Z}^{2}_{p}$, and all clients jointly generate random matrices $T_{i}\leftarrow\mathbb{Z}_{p}^{2\times 2}$ such that $\sum_{i\in[n]}T_{i}=0$. $s_{i}$ and $T_{i}$ forms the secret key held by each client. This interactive procedure allows clients to locally construct their own keys without relying on a trusted third party.
The decentralized architecture of DMCFE brings forth additional challenges posed by potentially dishonest clients, prompting the proposal of verifiable functional encryption (VFE) [13] as a solution. Nguyen et al.[14] proposed a verifiable decentralized multi-client functional encryption scheme for inner products (VDMCFE), along with its formal definitions and security proofs. This work extended the DMCFE scheme by constructing the Class Group-Based One-Time Decentralized Sum (ODSUM) and Label-Supporting DSUM (LDSUM) schemes, leveraging the Combine-then-Descend technique to improve the efficiency of correctness proofs for functional decryption key shares. Additionally, it incorporated zero-knowledge proofs such as range proofs to ensure the correctness of ciphertexts and the validity of the encrypted content.
@@ II-B Federated Learning
Biscotti[15] DP+SS
SecProbe[16] DP
ESFL[17] DP
SecureFL[18] PLHE+SS
AegisFL[6] HE
RVPFL[19] HE
FheFL[7] FHE
FEFL[20] 2DMCFE
BSR-FL[21] NIFE
@@
Note: Self-contained means the scheme works in a single-server setup without trusted third parties. Verifiability ensures the client follows protocol specifications. Fidelity [22] means no accuracy loss when no attack exists.
@@
  • FedAvg[1]. FedAvg is the most basic aggregation rule constructed when federated learning was first proposed. It directly computes the weighted sum of all local model to obtain the global model, where the weight of each client is determined by the proportion of its local dataset size to the total dataset size. However, FedAvg is not resistant to malicious client attacks, where the adversary manipulates the global model at will through submitted local models.
  • Krum[23]. This AGR selects global model update based on the squared distance score. Formally, for client $i$, the score is defined as $s(i)=\sum_{i\to j,n-f-2}\|W_{i}-W_{j}\|^{2}$ where $f$ represents the maximum number of malicious clients that can be tolerated and $i\to j$ denotes that $W_{j}$ belongs to the $n-f-2$ closest updates to $W_{i}$. Subsequently, the model with the lowest score or the average of several models with the lowest scores was chosen as the global model.
  • FLTrust[22]. The method requires the server to maintain a small clean dataset, which is used to train a trusted model update $g_{0}$. A trust score for each local model update $g_{i}$ is then computed by evaluating the cosine similarity between $g_{i}$ and $g_{0}$, followed by applying a ReLU operation. The score is used as the aggregation weight of the local model in the aggregation process.
  • AGR in FheFL[7]. In this schemes, the aggregation methods use the Euclidean distance between the local model $W^{t}_{i}$ and the global model $W^{t-1}$ of the previous round as a discriminator of the quality of the local model $W^{t}_{i}$.
  • AGR in BSR-FL[21]. This aggregation method constructs a qualified aggregation model using secure cosine similarity and an incentive-based Byzantine robustness strategy. In BSR-FL, the model from the $(t-1)$-th iteration $W^{t-1}$, serves as the baseline model. The trust score $cs^{t}_{i}$ for each local model $W_{i}^{t}$ is computed as the cosine similarity between $W_{i}^{t}$ and the baseline $W^{t-1}$, and then transformed using a parameterized sigmoid function centered at 0.5. However, the BSR does not verify the length of local models, leaving the potential for spoofing.
Existing aggregation rules are either not robust enough to resist adaptive attacks[22], or too complex to be applicable for fast verification of zero-knowledge proofs. Therefore, we propose a new aggregation rule that is applicable to zero-knowledge proofs and provides an effective defense against attacks such as adaptive attacks.
@@
    Differential Privacy Based FL. The Biscotti[15] combined differential privacy, Shamir secret shares (SS) and Multi-Krum aggregation rule to realize a private and secure federated learning scheme based on the blockchain. SecProbe[16] aggregated local models based on the computed utility scores and applies a functional mechanism to perturb the objective function to achieve privacy. Miao et al.[17] proposed an efficient and secure federated learning (ESFL) framework based on DP. Since the nature of differential privacy techniques is to hide the sensitive information of an individual by introducing random noise in the data or computational results, it cannot avoid affecting the accuracy of the model.
    Homomorphic Encryption Based FL. SecureFL[18] implemented FLTrust construction under encryption based on packed linear homomorphic encryption. However, in this scheme, the client distributes the local gradient to two servers through secret sharing, and the privacy protection of the gradient must rely on two non-colluding servers. AegisFL[6] improved the computational efficiency of multiple aggregation rules under ciphertexts by optimizing the homomorphic encryption scheme, but it also fails to eliminate dependency on two non-colluding servers. The RVPFL[19] scheme combined homomorphic encryption and Diffie-Hellman key agreement to achieve privacy protection and robustness detection, but relies on additional participants. FheFL[7] built the scheme on homomorphic encryption in a single server. However, a trusted key distribution process is necessary to ensure the secure generation of pairwise keys among users. The scheme does not validate the client behavior, and the malicious clients can easily destroy the decryption through wrong keys. Such schemes often use homomorphic encryption for privacy, computing aggregation on ciphertexts. Decryption typically requires a trusted party or two non-colluding servers. Shared decryption keys risk exposing honest clients’ data to key holders through collusion or eavesdropping.
    Functional Encryption Based FL. Chang et al.[20] constructed a privacy-preserving federated learning scheme FEFL via functional encryption. Qian et al.[24] designed another privacy-preserving federated learning scheme by combining functional encryption with multiple cryptographic primitives. However, both schemes do not address the detection of malicious clients. The BSR-FL scheme [21] combined functional encryption and blockchain technology to construct an efficient Byzantine robust privacy-preserving federated learning framework. However, the scheme still requires two non-colluding entities (the task publisher and the server) and does not validate the magnitude of the local models submitted by clients.
@@ III-A System Model
$lr$ local learning rate
@@ III-B Design Goals
    Byzantine Robustness. VFEFL can resist Byzantine attacks such as Gaussian attack[25], the adaptive attack[22] and Label Flipping attack, and obtain a high-quality global model.
@@ III-C Security Model
The protocol $\pi$ securely realizes a function $f$ if, for any adversary $\mathcal{A}$, there exists no environment $\mathcal{E}$ can distinguish whether it is interacting with $\pi$ and $\mathcal{A}$ or with the ideal process for $f$ and ideal adversary $\mathcal{A}^{*}$ with non-negligible probability. Formally, let $\text{REAL}_{\mathcal{A}}^{\mathcal{\pi}}(\lambda,\hat{x})$ and $\text{IDEAL}_{\mathcal{A}^{*}}^{f}(\lambda,\hat{x})$ denote the interaction views of $\mathcal{E}$ in the real-world and ideal-world models, respectively. For any $\mathcal{E}$, if the following holds:
@@ IV-B Zero Knowledge Proof
Notably, an interactive proof-of-knowledge protocol can be transformed into a non-interactive proof-of-knowledge through the application of the Fiat-Shamir heuristic [29].
@@ IV-C Verifiable Functional Encryption
Unlike traditional public key encryption, where decryption is an all-ro-nothing affair, functional encryption[30] can finely controls the amount of information that is revealed by a ciphertext to a given receiver. To capture adversarial correctness based on fine-grained control, verifiable functional encryption[13] was defined. Informally speaking, DVFE requires that, for any ciphertext $C$ passing a public verification procedure, there exists a unique message $m$ such that, for any valid function description $f$ and function key $dk_{f}$ passing public verification, the decryption algorithm, when given $C$, $dk_{f}$, and $f$ as inputs, outputs $f(m)$.
We follow the formal definition of VMCFE proposed in [14] to redefine our DVFE, including eight algorithms as following definition.
@@
  • KeyGen($pp$) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: Each client $\mathcal{C}_{i}$ generates own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs the verifiation key for ciphertexts $vk_{CT}$ and for functional key share $vk_{Dk}$, and the public key $pk$.
  • Encrypt$(ek_{i},x_{i},x_{0},\ell,\left(\ell_{Enc,j}\right)_{j\in[m]},pp)\to(C_{\ell,i},\pi_{Encrypt,i})$ : Takes as input an encryption key $ek_{i}$, a vector $x_{i}$ to encrypt, a vector $x_{0}$ used in the ciphertext proof protocol, a label $\ell$ and a set encryption labels $\left(\ell_{Enc,j}\right)_{j\in[m]}$ that ensures the ciphertexts of different dimensions in each round cannot be illegally aggregated. Outputs the ciphertext $C_{\ell,i}$ and a zero knowledge proof $\pi_{Encrypt,i}$.
  • VerifyCT($(C_{\ell,i})_{i\in[n]},\pi_{Encrypt,i},vk_{CT}$) $\to 1\text{ or }0$: Takes as input the ciphertext $C_{\ell,i}$ of a client, a verification key $vk_{CT}$. Outputs $1$ for accepting or 00 with a set of malicious clients $\mathcal{CC}_{CT}\neq\emptyset$ for rejecting.
@@
  • VerifyDK($(dk_{i,y})_{i\in[n]},\pi_{DKeyGenShare,i},vk_{DK},pp$) $\to 1\text{ or }0$: Takes as input functional decryption key shares $(dk_{i,y})_{i\in[n]}$, a verification key $vk_{DK}$. Outputs $1$ for accepting or 00 with a set of malicious clients $\mathcal{CC}_{DK}\neq\emptyset$ for rejecting.
  • DKeyComb($(dk_{i,y})_{i\in[n]},\ell_{y},pk$) $\to h^{dk_{y}}$: Takes as input all of the functional decryption key shares $(dk_{i,y})_{i\in[n]}$, a function label $\ell_{y}$ and public key $pk$. Outputs the functional decryption key $h^{dk_{y}}$.
@@ IV-D Our New Aggregation Rule
Inspired by the aggregation rule of FLTrust[22], we similarly introduce root dataset $D_{0}$ to bootstrap trust to achieve Byzantine robustness against malicious clients and consider both the direction and magnitudes of the local models of clients. However, FLTrust contains cosine similarity operations that are difficult to complete the verification using a simple zero-knowledge proof, which is avoided in our new aggregation rule.
@@ V-A Construction of DVFE
Although verifiable functional encryption schemes supporting weight aggregation for encrypting local models are now readily available, their application to privacy-preserving and robust federated learning faces the following two challenges. The first one is that the current ciphertext verification in verifiable functional encryption supports range proofs for individual elements but cannot verify Byzantine robustness rules. Secondly, the construction of zero-knowledge proofs for the Byzantine robust aggregation rule faces compatibility issues: existing inner-product arguments [31] and functional encryption schemes cannot be directly combined for verification. Specifically, the current inner-product argument fails to achieve the zero-knowledge property, and the ciphertext structure in functional encryption does not support direct verification of Byzantine robust aggregation rules.
@@
  • SetUp($\lambda$) $\to pp$: Takes as input the security parameter $\lambda$ and outputs the public parameters $pp=\left(\mathcal{PG},\mathcal{DG},h_{p},\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}^{\prime}_{1},\mathcal{H},\ell_{D}\right)$, where $\mathcal{PG}:=(\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T},p,g,h,e)\overset{\$}{\leftarrow}\mathrm{\texttt{PGGen}}(\lambda)$ is a pairing group, $g$ the generator of $\mathbb{G}_{1}$, $h$ the generator of $\mathbb{G}_{2}$, $\mathcal{DG}=(\tilde{s},f,\hat{h}_{p},\hat{G},F,\hat{G}^{p})\leftarrow\texttt{GenClassGroup}(\lambda,p)$ is a DDH group with an easy DL subgroup, $\hat{t}\leftarrow\mathcal{D}_{p}$ is sampled, and $h_{p}=\hat{h}_{p}^{\hat{t}}$ is set. $\mathcal{H}_{1}:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}^{2}$, $\mathcal{H}_{2}:\{0,1\}^{*}\rightarrow\mathbb{G}_{2}^{2}$, $\mathcal{H}^{\prime}_{1}:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}$, :$\mathcal{H}:\{0,1\}^{*}\rightarrow\mathbb{Z}_{p}$ are four full-domain hash function, and $\ell_{D}:=(\{0,1\}^{*})$ is the initialization label.
  • KeyGen($pp$) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: It is executed by each client $\mathcal{C}_{i}$ to generate own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs a verifiation key for ciphertexts $vk_{CT}$ and functional key share $vk_{Dk}$, and a public key $pk$. It takes $pp$ as input, and a client selects $\hat{k}_{i}=(\hat{k}_{i,1},\hat{k}_{i,2})\overset{\$}{\leftarrow}\mathbb{Z}_{p}^{2}$, $t_{i}=(t_{i,1},t_{i,2})\leftarrow\mathcal{D}^{2}_{p}$ and $s_{i}=(s_{i,1},s_{i,2})\overset{\$}{\leftarrow}\mathbb{Z}_{p}^{2}$, then computes $T_{i}=\left(h_{p}^{t_{i,b}}\right)_{b\in[2]}$ and
@@
  • Encrypt$(ek_{i},x_{i},x_{0},\ell,\left(\ell_{Enc,j}\right)_{j\in[m]},pp)\to(C_{\ell,i},\\ \pi_{Encrypt,i})$: Takes as input an encryption key $ek_{i}$, a vector $x_{i}$ to encrypt, a vector $x_{0}$ used in the ciphertext proof protocol, a label $\ell$ and a set encryption labels $\left(\ell_{Enc,j}\right)_{j\in[m]}$ that ensures the ciphertexts of different dimensions in each round cannot be illegally aggregated. The ciphertext is $C_{\ell,i}=\left(\left(u\right)^{s_{i}}\cdot(w_{j})^{x_{i,j}}\right)_{j\in[m]}$ where $u=\mathcal{H}_{1}(l)\in\mathbb{G}_{1}^{2}$, $w_{j}=\mathcal{H}_{1}^{{}^{\prime}}(\ell_{Enc,j})\in\mathbb{G}_{1}$. It re-computes $\mathrm{com}_{i}$ and $V=\prod_{j\in[m]}C_{\ell,i,j}$, then generates a proof $\pi_{Encrypt,i}$ for the relation $\mathcal{R}_{Encrypt}$ on input $\left(V,\mathrm{com}_{i},\ell,\ell_{Enc},y,x_{0}\right)$. The $\pi_{Encrypt,i}$ is described as
@@
  • VerifyCT($(C_{\ell,i})_{i\in[n]},\pi_{Encrypt,i},vk_{CT}$) $\to 1\text{ or }0$: Takes as input the ciphertext of a client $C_{\ell,i}$, proof $\pi_{Encrypt,i}$ and a verification key $vk_{CT}$. For $i\in[n]$: it verifies $\pi_{Encrypt,i}$ for the relation $\mathcal{R}_{Encrypt}$ and outputs $1$ for accepting or 00 with a set of malicious clients $\mathcal{CC}_{CT}\neq\emptyset$ for rejecting.
  • DKeyGenShare$(sk_{i},pk,y_{i},\ell_{y},pp)\to(dk_{y,i},\\ \pi_{DKeyGenShare,i})$: Takes as input a client secret key $sk_{i}=(s_{i},\ \hat{k}_{i},\ t_{i})$, public key $pk$, the function vector $y_{i}$ and a function label $\ell_{y}$. It computes $dk_{i}$ as
@@
  • VerifyDK($(dk_{i,y})_{i\in[n]},\pi_{DKeyGenShare,i},vk_{DK},pp$) $\to 1\text{ or }0$: Takes as input functional decryption key shares $(dk_{i,y})_{i\in[n]}$, proof $\pi_{DKeyGenShare,i}$ and a verification key $vk_{DK}$. For $i\in[n]$: it verifies $\pi_{DKeyGenShare,i}$ for the relation $\mathcal{R}_{DKeyGenShare}$ and outputs $1$ for accepting or 00 with a set of malicious clients $\mathcal{CC}_{DK}\neq\emptyset$ for rejecting.
  • DKeyComb($(dk_{i,y})_{i\in[n]},\ell_{y},pk$) $\to h^{dk_{y}}$: Takes as input all of the functional decryption key shares $(dk_{i,y})_{i\in[n]}$, a function label $\ell_{y}$ and public key $pk$. It computes $f^{d}=\left(\prod_{i\in[n]}d_{i,b}\right)_{b\in[2]}$ and solves the discrete logarithm problem on $\mathcal{DG}$ to obtain $d$, then computes $h^{dk_{y}}$ as
@@
$\mathrm{Adv_{DVFE}^{sta-ind}}(t,q_{E},q_{K})\leq$ $q_{E}\mathrm{Adv}^{\mathrm{zk}}_{\prod_{Encrypt}}(t)$

### v3

# Privacy-Preserving Federated Learning against Malicious Clients Based on Verifiable Functional Encryption

@@ I Introduction
Federated learning[FL2017] is an advanced and promising distributed learning paradigm that enables multiple clients to collaboratively train models without sharing local data. It embodies the principles of focused data collection and minimization, thereby mitigating many of the systemic privacy risks and costs inherent in traditional, centralized machine learning approaches. With the enactment of privacy protection and data security regulations in various countries, as well as the increase in public privacy awareness, the application of federated learning has gradually expanded e. g., healthcare, finance, and edge computing.
Although the federated learning framework offers a degree of privacy protection for client data during training, it brings new privacy and security issues [OpenFL]. One primary concern is the potential for privacy leakage through the local models uploaded by clients. With the advent and ongoing advancement of attack techniques, such as model inversion attacks[GIA, MIA, GIA2], submitting local models in plaintext form is no longer regarded as a secure method. Accordingly, obfuscation and encryption mechanisms have been adopted to address this issue. Another threat is attacks from malicious clients. Due to its distributed architecture, federated learning is susceptible to adversarial manipulation by malicious clients. A malicious client may make the accuracy of the global model drop rapidly by uploading malicious models, or destroy the decryption by uploading a malicious decryption key or a wrong ciphertext. The above issues are important for federated learning systems to achieve their core goal (training high quality models while preserving data privacy). Privacy-preserving mechanisms (such as encryption schemes) can potentially lead to difficulties in detection and verification, while detecting and verifying encrypted models of the clients may compromise specific information. Existing solutions [AegisFL, FheFL] cannot be deployed in the basic federated learning architecture (one server and multiple clients), and need help from two non-colluding servers or additional trusted third parties, which introduces additional limitations and consumption.
@@
    We employ a verifiable functional encryption scheme to encrypt local models in the federated learning, ensuring data privacy and correctness during encryption and decryption. The scheme supports verifiable evaluation of relations over multidimensional ciphertexts, making it well suited for federated learning scenarios. Furthermore, we formalize its definition and security model, and reduce its security in the proposed model to well-established computational hardness assumptions.
@@ II-A Functional Encryption
Functional encryption (FE) [FE_def, FE] is a paradigm supports selective computation on encrypted data, while traditional public key encryption only implements all-or-nothing decryption method. An authority in a FE scheme can generate restricted decryption keys to allow receiver to learn specific functions of the encrypted messages and nothing else. Goldwasser[MIFE2014] first introduced multi-input functional encryption (MIFE), which is a generalization of functional encryption to the setting of multi-input functions. Michel Abdalla [MIFE2017] constructed the first MIFE for a non-trivial functionality with polynomial security loss under falsifiable assumptions for a super-constant number of slots.
However, whereas the input can be large, the original definition of FE limits the source of the encrypted data to a single client, which is not compatible with a wide range of real-world scenarios where data needs to be collected from multiple sources. Multi-client functional encryption (MCFE)[DMCFE] was proposed to addresses this issue. Further, decentralized multi-client functional encryption (DMCFE)[DMCFE] eliminated reliance on a trusted party. In the DMCFE scheme instantiated over an asymmetric pairing group, each client independently samples $s_{i}\leftarrow\mathbb{Z}^{2}_{p}$, and all clients jointly generate random matrices $T_{i}\leftarrow\mathbb{Z}_{p}^{2\times 2}$ such that $\sum_{i\in[n]}T_{i}=0$. $s_{i}$ and $T_{i}$ forms the secret key held by each client. This interactive procedure allows clients to locally construct their own keys without relying on a trusted third party.
The decentralized architecture of DMCFE brings forth additional challenges posed by potentially dishonest clients, prompting the proposal of verifiable functional encryption (VFE) [VFE] as a solution. Nguyen et al.[VMCFE] proposed a verifiable decentralized multi-client functional encryption scheme for inner products (VDMCFE), along with its formal definitions and security proofs. This work extended the DMCFE scheme by constructing the Class Group-Based One-Time Decentralized Sum (ODSUM) and Label-Supporting DSUM (LDSUM) schemes, leveraging the Combine-then-Descend technique to improve the efficiency of correctness proofs for functional decryption key shares. Additionally, it incorporated zero-knowledge proofs such as range proofs to ensure the correctness of ciphertexts and the validity of the encrypted content.
@@ II-B Federated Learning
Biscotti[Biscotti] DP+$S$
SecProbe[SecProbe] DP
ESFL[ESFL] DP
SecureFL[SecureFL] PLHE+$S$
AegisFL[AegisFL] HE
RVPFL[RVPFL] HE
FheFL[FheFL] FHE
FEFL[FEFL] 2DMCFE
BSR-FL[BSRFL] NIFE
@@
Note: Self-contained means the scheme works in a single-server setup without trusted third parties. Verifiability ensures the client follows protocol specifications. Fidelity [FLtrust] means no accuracy loss when no attack exists.
@@
  • FedAvg[FL2017]. FedAvg is the most basic aggregation rule constructed when federated learning was first proposed. It directly computes the weighted sum of all local model to obtain the global model, where the weight of each client is determined by the proportion of its local dataset size to the total dataset size. However, FedAvg is not resistant to malicious client attacks, where the adversary manipulates the global model at will through submitted local models.
  • Krum[Krum]. This AGR selects global model update based on the squared distance score. Formally, for client $i$, the score is defined as $s(i)=\sum_{i\to j,n-f-2}\|W_{i}-W_{j}\|^{2}$ where $f$ represents the maximum number of malicious clients that can be tolerated and $i\to j$ denotes that $W_{j}$ belongs to the nf2n-f-2 closest updates to $W_{i}$. Subsequently, the model with the lowest score or the average of several models with the lowest scores was chosen as the global model.
  • FLTrust[FLtrust]. The method requires the server to maintain a small clean dataset, which is used to train a trusted model update $g_{0}$. A trust score for each local model update $g_{i}$ is then computed by evaluating the cosine similarity between $g_{i}$ and $g_{0}$, followed by applying a ReLU operation. The score is used as the aggregation weight of the local model in the aggregation process.
  • AGR in FheFL[FheFL]. In this schemes, the aggregation methods use the Euclidean distance between the local model $W^{t}_{i}$ and the global model $W^{t-1}$ of the previous round as a discriminator of the quality of the local model $W^{t}_{i}$.
  • AGR in BSR-FL[BSRFL]. This aggregation method constructs a qualified aggregation model using secure cosine similarity and an incentive-based Byzantine robustness strategy. In BSR-FL, the model from the (t1)(t-1)-th iteration $W^{t-1}$, serves as the baseline model. The trust score $cs^{t}_{i}$ for each local model $W_{i}^{t}$ is computed as the cosine similarity between $W_{i}^{t}$ and the baseline $W^{t-1}$, and then transformed using a parameterized sigmoid function centered at 0.5. However, the BSR does not verify the length of local models, leaving the potential for spoofing.
Existing aggregation rules are either not robust enough to resist adaptive attacks[FLtrust], or too complex to be applicable for fast verification of zero-knowledge proofs. Therefore, we propose a new aggregation rule that is applicable to zero-knowledge proofs and provides an effective defense against attacks such as adaptive attacks.
@@
    Differential Privacy Based FL. The Biscotti[Biscotti] combined differential privacy, Shamir secret shares (SS) and Multi-Krum aggregation rule to realize a private and secure federated learning scheme based on the blockchain. SecProbe[SecProbe] aggregated local models based on the computed utility scores and applies a functional mechanism to perturb the objective function to achieve privacy. Miao et al.[ESFL] proposed an efficient and secure federated learning (ESFL) framework based on DP. Since the nature of differential privacy techniques is to hide the sensitive information of an individual by introducing random noise in the data or computational results, it cannot avoid affecting the accuracy of the model.
    Homomorphic Encryption Based FL. SecureFL[SecureFL] implemented FLTrust construction under encryption based on packed linear homomorphic encryption. However, in this scheme, the client distributes the local gradient to two servers through secret sharing, and the privacy protection of the gradient must rely on two non-colluding servers. AegisFL[AegisFL] improved the computational efficiency of multiple aggregation rules under ciphertexts by optimizing the homomorphic encryption scheme, but it also fails to eliminate dependency on two non-colluding servers. The RVPFL[RVPFL] scheme combined homomorphic encryption and Diffie-Hellman key agreement to achieve privacy protection and robustness detection, but relies on additional participants. FheFL[FheFL] built the scheme on homomorphic encryption in a single server. However, a trusted key distribution process is necessary to ensure the secure generation of pairwise keys among users. The scheme does not validate the client behavior, and the malicious clients can easily destroy the decryption through wrong keys. Such schemes often use homomorphic encryption for privacy, computing aggregation on ciphertexts. Decryption typically requires a trusted party or two non-colluding servers. Shared decryption keys risk exposing honest clients’ data to key holders through collusion or eavesdropping.
    Functional Encryption Based FL. Chang et al.[FEFL] constructed a privacy-preserving federated learning scheme FEFL via functional encryption. Qian et al.[DMCFEFL] designed another privacy-preserving federated learning scheme by combining functional encryption with multiple cryptographic primitives. However, both schemes do not address the detection of malicious clients. The BSR-FL scheme [BSRFL] combined functional encryption and blockchain technology to construct an efficient Byzantine robust privacy-preserving federated learning framework. However, the scheme still requires two non-colluding entities (the task publisher and the server) and does not validate the magnitude of the local models submitted by clients.
@@ III-A System Model
lrlr local learning rate
@@ III-B Design Goals
    Byzantine Robustness. VFEFL can resist Byzantine attacks such as Gaussian attack[fang2020local], the adaptive attack[FLtrust] and Label Flipping attack, and obtain a high-quality global model.
@@ III-C Security Model
where $\overset{c}{\equiv}$ denotes computational indistinguishability, $\lambda$ represents the security parameter, and $\hat{x}$ includes all inputs.
@@ IV-B Zero Knowledge Proof
Notably, an interactive proof-of-knowledge protocol can be transformed into a non-interactive proof-of-knowledge through the application of the Fiat-Shamir heuristic [FiatShamir].
@@ IV-C Verifiable Functional Encryption
Unlike traditional public key encryption, where decryption is an all-ro-nothing affair, functional encryption[FEIP] can finely controls the amount of information that is revealed by a ciphertext to a given receiver. To capture adversarial correctness based on fine-grained control, verifiable functional encryption[VFE] was defined. Informally speaking, DVFE requires that, for any ciphertext $C$ passing a public verification procedure, there exists a unique message $m$ such that, for any valid function description $f$ and function key $dk_{f}$ passing public verification, the decryption algorithm, when given $C$, $dk_{f}$, and $f$ as inputs, outputs $f(m)$.
We follow the formal definition of VMCFE proposed in [VMCFE] to redefine our DVFE, including eight algorithms as following definition.
@@
  • KeyGen(pppp) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: Each client $\mathcal{C}_{i}$ generates own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs the verifiation key for ciphertexts $vk_{CT}$ and for functional key share $vk_{Dk}$, and the public key $pk$.
[…]
@@
[…]
@@ IV-D Our New Aggregation Rule
Inspired by the aggregation rule of FLTrust[FLtrust], we similarly introduce root dataset $D_{0}$ to bootstrap trust to achieve Byzantine robustness against malicious clients and consider both the direction and magnitudes of the local models of clients. However, FLTrust contains cosine similarity operations that are difficult to complete the verification using a simple zero-knowledge proof, which is avoided in our new aggregation rule.
@@ V-A Construction of DVFE
Although verifiable functional encryption schemes supporting weight aggregation for encrypting local models are now readily available, their application to privacy-preserving and robust federated learning faces the following two challenges. The first one is that the current ciphertext verification in verifiable functional encryption supports range proofs for individual elements but cannot verify Byzantine robustness rules. Secondly, the construction of zero-knowledge proofs for the Byzantine robust aggregation rule faces compatibility issues: existing inner-product arguments [Bulletproofs] and functional encryption schemes cannot be directly combined for verification. Specifically, the current inner-product argument fails to achieve the zero-knowledge property, and the ciphertext structure in functional encryption does not support direct verification of Byzantine robust aggregation rules.
@@
  • SetUp($\lambda$) $\to pp$: Takes as input the security parameter $\lambda$ and outputs the public parameters $pp=\left(\mathcal{PG},\mathcal{DG},h_{p},\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}^{\prime}_{1},\mathcal{H},\ell_{D}\right)$, where $\mathcal{PG}:=(\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T},p,g,h,e)\overset{\mathdollar}{\leftarrow}\mathrm{\texttt{PGGen}}(\lambda)$ is a pairing group, $g$ the generator of $\mathbb{G}_{1}$, $h$ the generator of $\mathbb{G}_{2}$, $\mathcal{DG}=(\tilde{s},f,\hat{h}_{p},\hat{G},F,\hat{G}^{p})\leftarrow\texttt{GenClassGroup}(\lambda,p)$ is a DDH group with an easy DL subgroup, $\hat{t}\leftarrow\mathcal{D}_{p}$ is sampled, and $h_{p}=\hat{h}_{p}^{\hat{t}}$ is set. $\mathcal{H}_{1}:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}^{2}$, $\mathcal{H}_{2}:\{0,1\}^{*}\rightarrow\mathbb{G}_{2}^{2}$, $\mathcal{H}^{\prime}_{1}:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}$, $\mathcal{H}:\{0,1\}^{*}\rightarrow\mathbb{Z}_{p}$ are four full-domain hash function, and $\ell_{D}:=(\{0,1\}^{*})$ is the initialization label.
  • KeyGen(pppp) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: It is executed by each client $\mathcal{C}_{i}$ to generate own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs a verifiation key for ciphertexts $vk_{CT}$ and functional key share $vk_{Dk}$, and a public key $pk$. It takes $pp$ as input, and a client selects $\hat{k}_{i}=(\hat{k}_{i,1},\hat{k}_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, $t_{i}=(t_{i,1},t_{i,2})\leftarrow\mathcal{D}^{2}_{p}$ and $s_{i}=(s_{i,1},s_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, then computes $T_{i}=\left(h_{p}^{t_{i,b}}\right)_{b\in[2]}$ and
@@
[…]
@@
[…]
@@
[…]
@@
$\mathrm{Adv_{DVFE}^{sta-ind}}(t,q_{E},q_{K})\leq\$$q_{E}\mathrm{Adv}^{\mathrm{zk}}_{\prod_{Encrypt}}(t)$